OpenStudy (anonymous):

Currently, Sam's coach is twice as old as Sam. When Sam started soccer 10 years ago, his coach was four times his age. How old is Sam now?

OpenStudy (tkhunny):

Start be defining things you do not know. C = Current Age of Sam's Coach. You do the other one. Then, start making words into equations.

OpenStudy (pawanyadav):

Let Sam age be x now Then his coach age be 2x now 10 years ago According to question, 2x-10=4(x-10) So x=15 which is current age of Sam.

OpenStudy (tkhunny):

{sigh} I guess we'll never know if @edemko123 can do it.